OpenStudy (anonymous):

FACTORING: 314x^3+1256x^2-7536 in order to solve using discriminate formula(b^2-4ac)?

OpenStudy (anonymous):

you sure you dont mean 314x^2+1256x-7536 ?

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

well 314x^3+1256x^2 +0x -7536 a x^3 +b x^2 + cx +d x^2 (a x+b)+d maybe that helps

OpenStudy (anonymous):

x^2 (314 x+1256)-7536

OpenStudy (anonymous):

could help.

OpenStudy (anonymous):

That's confusing, I'm just gonna give up, find another method

OpenStudy (anonymous):

it seems fishy cuzz (b^2-4ac) applies to second degree polynomials. Not 3rd. Unless i am missing something.

OpenStudy (anonymous):

then maybe it should be 314x^2+1256x-7536, even though I was pretty sure I combined them right

OpenStudy (anonymous):

lol maybe. the discriminant for a that cubic function in standard form is b^2 c^2-4 a c^3-4 b^3 d+18 a b c d-27 a^2 d^2 :O

OpenStudy (anonymous):

or since you got no coeeficient in the x term -d (27 a^2 d+4 b^3)

OpenStudy (anonymous):

I'm not sure that's the result I'm looking for

OpenStudy (anonymous):

314 (x-2) (x^2+6 x+12) ( this ) ooh ^ now you can do discriminant :D

OpenStudy (anonymous):

ohh that's factored? or...

OpenStudy (anonymous):

that is 314x^3+1256x^2-7536 factored to 314 (x-2) (x^2+6 x+12) You can apply the discriminant to (x^2+6 x+12) but idk what for lol

OpenStudy (anonymous):

I was told to use the discriminant to find out the zeros by fundamental theorem?

OpenStudy (anonymous):

Can you give the original problem in the way it was originally stated?

OpenStudy (anonymous):

Option 1 - Cylinder You will need the following materials to find the volume of a cylinder: A cylindrical object such as a soup can or thermos Ruler or tape measure Graphing technology (e.g., graphing calculator or GeoGebra) Procedure: Measure and record the diameter and height of the cylindrical object you have chosen in inches. Round to the nearest whole number. Apply the formula of a right circular cylinder (V = r2h) to find the volume of the object. (Note: Be sure to find the radius from the diameter measurement by dividing by 2.) Now suppose you knew the volume of this object and the relation of the height to the radius, but did not know the radius. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the radius. Rewrite the formula using the variable x for the radius. Substitute the value of the volume found in step 2 for V and express the height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the radius, then the expression for the height will be (x + 4). Simplify the equation and write it in standard form. Multiply each term in the equation by 100 to eliminate any decimals, if necessary. Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem.

OpenStudy (anonymous):

314x^3+1256x^2-7536 First of all getting this factored is a pain in the butt. but since i gave it to you it's the following 314 (x-2) (x^2+6 x+12) ^ 314 dont matter ^ x-2=0 means solving for x will give you 2 as a root ^ (x^2+6 x+12) part... (x^2+6 x+12) you apply discriminant. a b c (b^2-4ac) plug in numbers 6^2-4(1)(12) 36-4*12 36-48=-12 is discriminant . If i remember right if D <0 then it has imaginary zeros.

OpenStudy (anonymous):

Okay what were the height, radius, and volume?

OpenStudy (anonymous):

h=6, d=4, r=2, v=75.36

OpenStudy (anonymous):

Okay, cool. So then, we're supposed to assume we don't know the radius or the height. We just know that r = h-4 and that volume = 75.36

OpenStudy (anonymous):

Does that much make sense to you?

OpenStudy (anonymous):

I thought the h=x+4 ?

OpenStudy (anonymous):

Yes. That's equivalent.

OpenStudy (anonymous):

yes and I after putting it all together and simplifying I got 314x^3+1256x^2-7536=0 ?

OpenStudy (anonymous):

Oh, you're taking out all of the decimals.

OpenStudy (anonymous):

yeah to make life easier

OpenStudy (anonymous):

Haha I tend to think that makes life a good bit harder. You know, since 314 is a much different number than 3.14.

OpenStudy (anonymous):

What are you suggesting?

OpenStudy (anonymous):

after working on this for two hours I'm sure my answer should come from 314(x-2)(x^2+6x+12) ?

OpenStudy (anonymous):

not if you made 3.14 into 314

OpenStudy (anonymous):

I see what she did. She just multiplied the whole equation by 100.

OpenStudy (anonymous):

is there a difference in the end? even the directions instructed that I get rid of all demcials

OpenStudy (anonymous):

Yeah, there's no problem with it. It was just confusing because I didn't realize what you were doing.

OpenStudy (anonymous):

Factor out a GCF of 314 to get: \(\Large 314(x^3 + 4x^2 + 24) = 0\)

OpenStudy (anonymous):

And that factors from there the way you had it. \(\Large 314(x-2)(x^2+6x+12) =0\) From the first factor, (x-2) you get what root?

OpenStudy (anonymous):

To get the roots that go along with a single factor, set the factor equal to 0 and solve for x.

OpenStudy (anonymous):

I wouldn't use discriminant at all??

OpenStudy (anonymous):

No. Once a polynomial is factored, you can find the solutions by setting each factor equal to 0. Let me show you an example. \(\Large (x+4)(x-1)(x^2+x+20) = 0\) x+4 =0 gives me x=-4 x-1 = 0 gives me x=1 \(\large x^2 + x +20 =0\) I need to use the quadratic formula for this one, and I get 2 imaginary solutions.